The mean playing time for a large collection of compact discs is 34 minutes, and the standard deviation is 4 minutes.

(a) What value (in minutes) is 1 standard deviation above the mean? One standard deviation below the mean? What values are 2 standard deviations away from the mean?

1 standard deviation above the mean

1 standard deviation below the mean

2 standard deviations above the mean

2 standard deviations below the mean

(b) Assuming that the distribution of times is mound shaped and approximately symmetric, approximately what percentage of times are between 26 and 42 minutes? (Hint: See Example 3.19. Use the Empirical Rule.)

%

Less than 22 min or greater than 46 min?

%

Less than 22 min?

%
Answer #1

a)

Mean =34

Std = 4

1 standard deviation above the mean: 34+4 = 38

1 standard deviation below the mean : 34-4 = 30

2 standard deviations above the mean: 34+8 = 42

2 standard deviations below the mean:34-8 = 26

b)

Assuming that the distribution of times is mound-shaped and approximately symmetric, approximately what percentage of times are between 26 and 42 minutes? (Hint: See Example 3.19. Use the Empirical Rule.)

95%

Less than 22 min or greater than 46 min?

100%-99.7% = 0.3%

Less than 22 min?

0.3/2%=0.15%